Circular No. 7696

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dz IN UGC 471 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles

(Coddenham, England) of an apparent supernova (mag 18.3) on eight unfiltered CCD frames (limiting mag 19.0–19.5) secured between Aug. 27.970 and 28.125 UT for the U.K. Nova/Supernova Patrol with a 0.36-m Schmidt-Cassegrain telescope. The new object is located at α = 0^{h}44^{m}43^{s}.15, δ = +45^{o}25′10′′.4 (equinox 2000.0), which is 13′′ east and 6′′ west of the center of UGC 471. In response to an appeal from the discoverer, M. Armstrong (Rolvenden, England) also imaged SN 2001dz on Aug. 28.031, and Boles found the star at mag 18.0 on another image taken on Aug. 28.940 (which produced position end figures 43^{s}.11, 10′′.6). SN 2001dz was absent from a master image taken by Boles on July 25 (limiting mag 19.0), and it is not present on second Palomar Sky Survey blue and red images from 1989.

SUPERNOVA 2001ea IN MCG +05-54-38 B. A. Beutler and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.2) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 28.4 and 29.4 UT. The new object is located at α = 23^{h}08^{m}04^{s}.68, δ = +28^{o}34′06′′.1 (equinox 2000.0), which is 10′′.3 east and 4′′.2 south of the nucleus of MCG +05-54-38. A KAIT image taken on Aug. 19.4 showed nothing at this position (limiting mag ∼ 19.0).

V4739 SAGITTARII K. Vanlandingham, Columbia University; R. M. Wagner, LBT Obser-

vatory; and S. G. Starrfield, Arizona State University, write: “An optical spectrum (range 325–900 nm; resolution 0.8 nm) was obtained of V4739 Sgr on Aug. 28.15 UT with the 2.4-m Hiltner telescope (+ CCDS spectrograph) of the MDM Observatory on Kitt Peak. The spectrum exhibits very broad emission lines of the Balmer series of hydrogen, He I, N II 568-nm, the C III/N III blend at 464–465 nm, and O I 777.4- and 844.6-nm superposed on a reddened continuum. Interstellar absorptions, particularly Na D at 589.3 nm, are prominent, suggesting that the object is moderately to heavily reddened. The FWZI and FWHM of Hα emission are 8044 and 5510 km/s, respectively, implying a mean expansion velocity of ∼ 2750 km/s. The equivalent widths of Hα and Hβ emission are 24 and 16 nm, respectively. The spectrum is typical of a He/N-class nova early in its evolution.”

2001 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green